OpenStudy (anonymous):

Find the value of the expression: -15 + (-10) + 25

OpenStudy (yrelhan4):

just like the previuos one.. -15 + (-10) =-15 - 10=-25 -25 + 25=0

OpenStudy (anonymous):

I'm confused .

OpenStudy (yrelhan4):

look a + (-b)=a-b got this part?

OpenStudy (hba):

Let me show you something (-)(-)=+ (-)(+)=+ (+)(+)=+ Got it @misswiss123

OpenStudy (hba):



OpenStudy (anonymous):

@hba - {(-)(-)=+ (-)(+)=+ (+)(+)=+} This is a bit confusing!

OpenStudy (anonymous):

a + (+b) = a+b a + (-b) = a-b a - (+b) = a-b a - (-b) = a+b

OpenStudy (anonymous):

so the answer would be?

OpenStudy (anonymous):

what do u think the answer would be?

OpenStudy (anonymous):

I don't know